When Benzoic acid is combusted, -3226.6 kJ/mol of heat is released. Calculate the heat of formation for benzoic acid. Show all steps.(Benzoic acid = C7H6O2 )

Thanks!!
Answer #1

Combustion reaction of Benzene is:

2 C6H5COOH + 15 O2 —> 14 CO2 + 6 H2O .

C6H5COOH + 15/2 O2 —> 7 CO2 + 3 H2O .  Hcomb = -3226.6 kJ

From literature

Hf(O2) = 0.0 kJ/mol

Hf (CO2) = -393.51 kJ/mol

Hf(H2O) = -285.8 kJ/mol

Hf( C6H5COOH ) = ?

From reaction

Hcomb = Hf( products) - Hf(reactants)

Hcomb = (3*Hf(H2O)) + (7*Hf (CO2)) - Hf( C6H5COOH ) - (15/2 * Hf(O2))

-3226.6 kJ/mol = (3*(-285.8)) + (7*(-393.51)) - Hf( C6H5COOH )

Hf( C6H5COOH ) = -3611.97-3226.6 = - 385.37 kJ/mol